# AgEval: A Benchmark for Zero-Shot and Few-Shot Plant Stress Phenotyping with Multimodal LLMs

### v1

@@ 3.2 Full (8)-shot Performance
The introduction of eight examples per task (8-shot learning) leads to significant changes in model performance (Figure 3). For identification tasks, GPT-4o demonstrates remarkable improvement, achieving the highest MRR of 0.81, a substantial increase from its zero-shot performance. This suggests that GPT-4o effectively leverages additional context to enhance its understanding of the identification tasks. In classification and quantification tasks, Claude-3.5-sonnet is the top performer with an MRR of 0.66. GPT-4o maintains strong performance in this category as well, with an MRR of 0.55. The consistent high performance of GPT-4o across both task types indicates its versatility in addressing various aspects of this benchmark.
Figure 3: 8-shot comparative performance of LLMs.
Notably, most models show improved performance with 8-shot learning compared to zero-shot, particularly in more complex classification and quantification tasks. This underscores the value of providing even a small number of examples in enhancing model accuracy. However, the degree of improvement varies across models, highlighting differences in their ability to utilize additional context effectively when tackling the diverse challenges presented.
Figure 5: Performance comparison on individual tasks of the AgLLM benchmark across different shot settings (0 to 8 shots) for top-4 performing LLMs. The plots showcase how each task’s performance evolves with increasing shot counts, revealing model-specific and task-specific trends in few-shot learning capabilities
Identification Tasks (F1 Score): In 8-shot identification, GPT-4o significantly improves, achieving an average F1 score of 73.37 and outperforming in 5 out of 7 tasks. It excels in Durum Wheat (95.94) and Bean Leaf Lesions (86.90) identification. Gemini-pro-1.5 maintains strong performance (62.72), leading in DeepWeeds identification (49.96). Interestingly, the Dangerous Insects task shows a slight performance decrease for some models, possibly due to the already high zero-shot performance.
Classification and Quantification Tasks (NMAE): For 8-shot classification and quantification, Claude-3.5-sonnet maintains the lead with an average NMAE of 12.78, excelling in 3 out of 5 tasks, including Yellow Rust 19 classification (16.04) and InsectCount quantification (5.75). Gemini-pro-1.5 shows substantial improvement (13.74), particularly in IDC classification (12.04). It’s worth noting that LLaVA’s performance unexpectedly deteriorates with increased shot numbers, often resulting in unparsable outputs and incorrect predictions. This anomaly suggests potential limitations in LLaVA’s in-context learning capabilities for the AgEval benchmark tasks. These results underscore the varying impacts of few-shot learning across models and tasks within AgEval, highlighting the importance of model selection based on specific agricultural task requirements and the potential influence of factors like image quality on performance.
@@ 3.3 Relevance of Examples in Few-Shot Learning
The relevance of examples in few-shot learning significantly influences model performance across various identification tasks. As shown in Table 3, the presence of exact category examples (Bullseye) consistently improves F1 scores, with an average increase of 15.38% across all shot settings. This impact is most pronounced in the 1-shot scenario (+21.66%), highlighting the role of relevant examples when context is limited. Conversely, the absence of exact category examples (Non-Bullseye) has a minimal overall impact (-0.12% on average), suggesting that models demonstrate robustness to irrelevant examples, especially as the number of shots increases.
The benefits of relevant examples persist across different shot counts, albeit with diminishing returns. The average Bullseye impact decreases from +21.66% in 1-shot to +14.70% in 8-shot settings, indicating that while relevant examples remain beneficial, their individual importance lessens with increased context. Dataset-specific variations are substantial: Soybean Diseases shows the highest Bullseye impact (+49.05% in 1-shot), while Bean Leaf Lesions exhibits more modest improvements (+11.46% in 1-shot). The Dangerous Insects dataset presents an exception, showing negative impacts even with Bullseye examples (-9.95% average impact), which may indicate domain-specific challenges. These findings underscore the importance of example selection in few-shot learning, particularly in low-shot scenarios, and suggest that the effectiveness of this approach varies across different domains within identification tasks. Please note that Llava was excluded from this analysis due to failure in few-shot learning for our benchmark (see Figure 5).
@@ 3.4 Intra-task Uniformity
Regarding datasets, Soybean Diseases exhibited the highest average CV (81.29%), indicating significant variability in model performance across its classes, potentially due to its low image resolution. Conversely, Durum Wheat showed the lowest average CV (14.28%), implying more uniform performance. DeepWeeds and Mango Leaf Disease also demonstrated high variability (CV >40%), highlighting challenges in consistently identifying diverse species and conditions.
These findings underscore the importance of considering not just overall accuracy, but also consistency across classes when selecting models for agricultural identification tasks. The substantial variations observed across datasets and models highlight the need for targeted improvements in LLM training and fine-tuning to enhance performance uniformity. Detailed plots for each dataset’s classes are provided in the supplement, offering deeper insights into class-specific performance variations.
@@ 3.5 Key Findings
(1) GPT-4o demonstrates superior adaptability across AgEval tasks, showing significant improvement with few-shot learning (F1 score increase from 46.24 to 73.37).
(2) Example relevance significantly impacts few-shot learning: On averge, exact category examples (bullseyes) improve F1 scores by 15.38% , while related examples from different classes have negligible impact.
(3) LLM performance within datasets differs widely: Coefficient of Variation ranges from 26.02% (GPT-4o) to 58.03% (Claude-3-haiku), indicating that even top-performing models may struggle with consistent accuracy across all classes in plant stress phenotyping tasks.
(4) Different models have their own strengths: Gemini-pro-1.5 excels in zero-shot identification (MRR 0.69), while GPT-4o leads in zero-shot classification/quantification (MRR 0.70). And Claude Sonnet-3.5 consistently performs better in classification/quantification tasks
@@ 4 Conclusions
This study introduces AgEval, a curated benchmark for evaluating state-of-the-art multimodal LLMs on plant stress phenotyping tasks. We assembled a diverse set of tasks from public and internal sources, focusing on identification, classification, and quantification across various crops and stress types. Our evaluation methodology examines zero-shot and few-shot performance, the impact of example relevance on predictions, and quantifies the variation in model performance across different classes within each task using the Coefficient of Variation. The results demonstrate the potential of multimodal LLMs in addressing plant stress phenotyping challenges, with varying strengths observed across different models and task categories. This work establishes a strong baseline benchmark for future LLMs to test their capabilities in agricultural contexts and lays the groundwork for more diverse benchmarks in this domain.
Future research could expand the scope to cover a broader range of agricultural tasks beyond plant stress phenotyping. Additionally, as our analysis was limited to 8 shots, future work could explore increased shot counts for tasks where performance did not plateau, to get further insights. Models could be fine-tuned to specifically improve performance on non-bullseye examples and enhance intra-task uniformity, potentially increasing consistency across different classes within each task. Finally, it will be important to assess the resource requirements of deploying these models in real-world agricultural settings (computational costs, continuous data updates, integration with existing agricultural systems, carbon footprint, as well as data privacy implications).
@@
Insect Count:
@@
PlantDoc (Disease Quantification)
Analyze this image of a leaf to get the total percentage of affected leaf. The images are of several plant leaf-like Apple Scab Leaf, Apple rust leaf, Bell_pepper leaf spot, Corn leaf blight, Potato leaf early blight, etc. The affected area is: diseased leaf area / total image area. Provide your answer in the following JSON format:{"prediction": "number"}Replace "number" with your best estimate of the percent on your analysis of the image. The number should be entered exactly as a whole number (without any symbols) in a range of {expected_classes} The response should start with { and contain only a JSON object (as specified above) and no other text.
@@
Table S1 provides a comprehensive overview of the datasets used in the AgEval benchmark. It categorizes each dataset based on its primary task (Identification, Classification, or Quantification) and subcategory (e.g., Seed Morphology, Foliar Stress, Pests). The table includes key information such as the number of images, classes, year of creation, geographical location, and the evaluation metric used. This diverse collection of datasets covers various aspects of plant stress phenotyping, ranging from seed quality assessment to disease severity classification across different crops and regions.
@@
Dataset Category Subcategory Description # of Images # of Classes Year Location Metric
Durum Wheat [40, 16] I Seed Morphology Wheat variety identification 333 3 2019 Turkey F1
Soybean Seeds [17] I Seed Morphology Soybean quality prediction 5513 5 N/A N/A F1
Mango Leaf Disease [18, 19] I Foliar Stress Mango leaf disease classification 4000 8 2022 Bangladesh F1
Bean Leaf Lesions [20] I Foliar Stress Bean leaf lesion type classification 1171 3 N/A N/A F1
Soybean Diseases [21] I Foliar Stress Soybean stress identification 25,000 9 2016 United States F1
Dangerous Insects [23] I Pests Harmful insects identification 1591 15 N/A N/A F1
DeepWeeds [41, 22] I Pests Weeds species identification 17,509 9 2019 Australia F1
Yellow Rust 19 [24, 25, 26] C Disease Severity Wheat yellow rust severity 20,421 6 2021 Turkey NMAE
FUSARIUM 22 [27, 28, 29] C Disease Severity Chickpea fusarium wilt severity 19,339 5 2023 Turkey NMAE
IDC [30] C Stress Tolerance Soybean stress severity 1,972 5 2015 United States NMAE
InsectCount [31] Q Pest Count Insect count in images 1204 - 2021-2022 N/A NMAE
PlantDoc [32, 33] Q Disease Percentage of the leaf that is diseased 1176 - N/A N/A NMAE
@@
Figure S2 provides a detailed examination of intra-task uniformity across different datasets in the AgEval benchmark. Each subfigure represents a specific dataset, showcasing the F1 scores for the highest, median, and lowest performing classes based on 0-shot performance. The visualization for each class displays both the 0-shot F1 score (solid bars) and the additional gain in F1 score achieved with 8-shot learning (hatched bars) for all six evaluated models. This comprehensive view highlights the significant performance disparities among classes within each task, supporting our finding that the coefficient of variance (CV) ranges from 26.02% to 58.03% across models. The stark differences between the highest and lowest performing classes underscore the need for subject matter expertise to achieve reliable performance, especially for ”difficult” classes.

### v2

@@ 3.2 Full (8)-shot Performance
The introduction of eight examples per task (8-shot learning) leads to significant improvements in model performance (Figure 3), achieving results that traditionally require thousands of annotated examples. For identification tasks, GPT-4o achieves the highest MRR of 0.81, a substantial increase from its zero-shot performance. In classification and quantification tasks, Claude-3.5-sonnet emerges as the top performer with an MRR of 0.66, while GPT-4o maintains strong performance with an MRR of 0.55.
Most models show improved performance with 8-shot learning compared to zero-shot, particularly in complex tasks, highlighting the efficiency of VLMs in learning from limited data.
Identification Tasks (F1 Score): In 8-shot identification, GPT-4o achieves an average F1 score of 73.37, outperforming in 5 out of 7 tasks. It excels in Durum Wheat (95.94) and Bean Leaf Lesions (86.90) identification. Gemini-pro-1.5 maintains strong performance (62.72), leading in DeepWeeds identification (49.96).
Classification and Quantification Tasks (NMAE): For 8-shot classification and quantification, Claude-3.5-sonnet leads with an average NMAE of 12.78, excelling in 3 out of 5 tasks. Gemini-pro-1.5 shows substantial improvement (13.74), particularly in IDC classification (12.04).
LLaVA’s performance unexpectedly deteriorates with increased shot numbers, suggesting potential limitations in its in-context learning capabilities for AgEval tasks. These results underscore the varying impacts of few-shot learning across models and tasks, highlighting the importance of model selection based on specific agricultural task requirements.
@@ 3.3 Relevance of Examples in Few-Shot Learning
The relevance of examples in few-shot learning significantly influences model performance across various identification tasks. As shown in Table 3, the presence of exact category examples (Bullseye) consistently improves F1 scores, with an average increase of 15.38% across all shot settings. This impact is most pronounced in the 1-shot scenario (+21.66%). The absence of exact category examples (Non-Bullseye) has a minimal overall impact (-0.12% on average), suggesting VLM robustness to less relevant examples, especially as the number of shots increases.
The benefits of relevant examples persist across different shot counts, albeit with diminishing returns. The average Bullseye impact decreases from +21.66% in 1-shot to +14.70% in 8-shot settings. Dataset-specific variations are substantial: Soybean Diseases shows the highest Bullseye impact (+49.05% in 1-shot), while Bean Leaf Lesions exhibits more modest improvements (+11.46% in 1-shot). The Dangerous Insects dataset presents an interesting case, showing slight negative impacts even with Bullseye examples (-9.95% average impact), which may indicate domain-specific nuances. These findings underscore the importance of example selection in few-shot learning, particularly in low-shot scenarios. Please note that Llava was excluded from this analysis due to challenges in few-shot learning for our benchmark (see Figure 5).
@@ 3.4 Intra-task Uniformity
Regarding datasets, Soybean Diseases exhibited the highest average CV (81.29%), indicating variability in model performance across its classes, potentially due to its low image resolution. Conversely, Durum Wheat showed the lowest average CV (14.28%), implying more uniform performance. DeepWeeds and Mango Leaf Disease also demonstrated higher variability (CV >40%), highlighting areas for targeted improvements in VLM training to enhance performance uniformity.
These findings underscore the importance of considering not just overall accuracy, but also consistency across classes when selecting models for agricultural identification tasks. Detailed plots for each dataset’s classes are provided in the supplement.
@@ 3.5 Key Findings
(1) GPT-4o demonstrates strong adaptability across AgEval tasks, showing significant improvement with few-shot learning (F1 score increase from 46.24 to 73.37).
(2) Example relevance significantly impacts few-shot learning: On average, exact category examples (bullseyes) improve F1 scores by 15.38%, while related examples from different classes have minimal impact.
(3) VLM performance within datasets shows some variation: Coefficient of Variation ranges from 26.02% (GPT-4o) to 58.03% (Claude-3-haiku), indicating opportunities for further refinement in achieving consistent accuracy across all classes in plant stress phenotyping tasks.
(4) Different models exhibit complementary strengths: Gemini-pro-1.5 excels in zero-shot identification (MRR 0.69), while GPT-4o leads in zero-shot classification/quantification (MRR 0.70). Claude Sonnet-3.5 consistently performs well in classification/quantification tasks, showcasing the diverse capabilities of VLMs in agricultural applications.
@@ 4 Conclusions
This study introduces AgEval, a benchmark for evaluating Vision Language Models (VLMs) on plant stress phenotyping tasks. We assembled diverse tasks across crops and stress types. Our evaluation examines zero-shot and few-shot performance, example relevance impact, and performance consistency using Coefficient of Variation. Results show VLMs’ potential in addressing plant stress phenotyping challenges, with complementary strengths across models and tasks. This work establishes a baseline for future VLMs in agricultural contexts.
VLMs demonstrate adaptability to specialized agricultural tasks, with improvements in few-shot learning. Intra-task uniformity variation highlights refinement opportunities. These findings show VLMs are scalable solutions for plant stress phenotyping with minimal context. While VLMs may not match specialized models’ peak performance, they offer valuable flexibility and reduced data requirements for practical agricultural applications
Future research could expand to broader agricultural tasks beyond plant stress phenotyping. Exploring increased shot counts could provide further insights. Fine-tuning models could enhance non-bullseye example performance and improve intra-task uniformity. Assessing practical deployment aspects in real-world settings, including computational requirements, data update strategies, integration with existing systems, and environmental and privacy considerations, will be crucial.
@@
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help(keyval) Package keyval Error: breakanywhere undefinedSee the keyval package documentation for explanation.I can’t help
@@
Insect Count: (keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help(keyval) Package keyval Error: breakanywhere undefinedSee the keyval package documentation for explanation.I can’t help
@@
PlantDoc (Disease Quantification) (keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help(keyval) Package keyval Error: breakanywhere undefinedSee the keyval package documentation for explanation.I can’t help
Analyze this image of a leaf to get the total percentage of affected leaf. The images are of several plant leaf-like Apple Scab Leaf, Apple rust leaf, Bell_pepper leaf spot, Corn leaf blight, Potato leaf early blight, etc. The affected area is: diseased leaf area / total image area. Provide your answer in the following JSON format:{"prediction": "number"}Replace "number" with your best estimate of the percent on your analysis of the image. The number should be entered exactly as a whole number (without any symbols) in a range of {expected_classes} The response should start with { and contain only a JSON object (as specified above) and no other text.only a JSON object (as specified above) and no other text.
@@
Table S1 provides a comprehensive overview of the datasets used in the AgEval benchmark. It categorizes each dataset based on its primary task (Identification, Classification, or Quantification) and subcategory (e.g., Seed Morphology, Foliar Stress, Pests). The table includes key information such as the number of images, classes, year of creation, geographical location, and the evaluation metric used. This diverse collection of datasets covers various aspects of plant stress phenotyping, ranging from seed quality assessment to disease severity classification across different crops and regions. Table S2 provides a comparison of the performance of traditional models on these datasets.
@@
Dataset Category Subcategory Description # of Classes Year Location Metric
Durum Wheat [41, 16] I Seed Morphology Wheat variety identification 3 2019 Turkey F1
Soybean Seeds [17] I Seed Morphology Soybean quality prediction 5 N/A N/A F1
Mango Leaf Disease [18, 19] I Foliar Stress Mango leaf disease classification 8 2022 Bangladesh F1
Bean Leaf Lesions [20] I Foliar Stress Bean leaf lesion type classification 3 N/A N/A F1
Soybean Diseases [21] I Foliar Stress Soybean stress identification 9 2016 United States F1
Dangerous Insects [23] I Pests Harmful insects identification 15 N/A N/A F1
DeepWeeds [42, 22] I Pests Weeds species identification 9 2019 Australia F1
Yellow Rust 19 [25, 26, 27] C Disease Severity Wheat yellow rust severity 6 2021 Turkey NMAE
FUSARIUM 22 [28, 29, 30] C Disease Severity Chickpea fusarium wilt severity 5 2023 Turkey NMAE
IDC [31] C Stress Tolerance Soybean stress severity 5 2015 United States NMAE
InsectCount [32] Q Pest Count Insect count in images - 2021-2022 N/A NMAE
PlantDoc [33, 34] Q Disease Percentage of the leaf that is diseased - N/A N/A NMAE
Table S2: Performance of the Traditional Models on Agricultural Image Datasets
Dataset Method/Approach Used Reported Metric (Score) Train Validation Test
Durum Wheat [41, 16] Transfer learning with EfficientNetB3 F1 Score (100) 227(70%) 49 (15%) 49 (15%)
Soybean Seeds [17] Transfer learning with ResNet50 Accuracy (89) 4410(80%) - 1103(20%)
Mango Leaf Disease [18, 19] Transfer learning with EfficientNetB3 Accuracy (100) 3200(80%) 480(12%) 320(8%)
Bean Leaf Lesions [20] Hybrid Model (ViT, SVM) F1 score (91) 974 (84%) 133 (11%) 60 (5%)
Soybean Diseases [21] Convolutional neural network Accuracy (94) 53266(81%) 5918(9%) 6576(10%)
Dangerous Insects [23] Transfer learning with Xception Accuracy (77) 1272 (80%) 287 (18%) 32 (2%)
DeepWeeds [42, 22] Transfer learning with ResNet26 F1 score (91) 11205(64%) 2801(16%) 3501(20%)
Yellow Rust 19 [25, 26, 27] CNN-CGLCM with SVM Accuracy (92) 13500 (90%) - 1500 (10%)
FUSARIUM 22 [28, 29, 30] Hybrid Classifier (ViT,CatBoost) F1 score (75) 2950(68%) 521 (12%) 868(20%)
IDC [31] Hierarchical classification Accuracy (96) 1479(75%) - 493 (25%)
InsectCount [32] Internal dataset No baseline published
PlantDoc [33, 34] No baseline exists on this data for this task
@@
Figure S2 provides a detailed examination of intra-task uniformity across different datasets in the AgEval benchmark. Each subfigure represents a specific dataset, showcasing the F1 scores for the highest, median, and lowest performing classes based on 0-shot performance. The visualization for each class displays both the 0-shot F1 score (solid bars) and the additional gain in F1 score achieved with 8-shot learning (hatched bars) for all six evaluated models. This comprehensive view highlights the significant performance disparities among classes within each task, supporting our finding that the coefficient of variance (CV) ranges from 26.02% to 58.03% across models. The stark differences between the highest and lowest performing classes underscore the need for subject matter expertise to achieve reliable performance, especially for "difficult" classes.